what is the perimeter in feet if a window is 20 yards wide by 936 feet high

Answer by checkley79(3341)

30 yards is 3*30 (yds.) = 90 feet.
Perimeter=2L+2W
Perimeter=2*936+2*90
Perimeter=1872+180=2052 feet ans.